Plot the following P and Q combination. (0,25),(1,20),(2,15),(3,10),(4,5) and using the equation of a line, what is algebraic formula of this demand curve?
1
Expert's answer
2021-08-23T13:19:00-0400

The general equation of a straight line is given by

"y=mx+c"

c=y intercept

m=slope of curve


From the given points we can plot the straight line and then we can find both c and m




"y=mx\\when\space x=0, y=c\\x=0,y=25\\\therefore c=25"


Coincides at point (1,200. Putting these values in the equation.

"y=mx+c\\20=m+25\\m=20-25\\slope=-5\\intercept=25"


hence the equation of the line is given by

"y=-5x+25"